### v1

@@ Abstract
We present a simultaneous sensor-based inspection and footprint coverage (SIFC) planning and control design with applications to autonomous robotic crack mapping and filling. The main challenge of the SIFC problem lies in the coupling of complete sensing (for mapping) and robotic footprint (for filling) coverage tasks. Initially, we assume known target information (e.g., crack) and employ classic cell decomposition methods to achieve complete sensing coverage of the workspace and complete robotic footprint coverage using the least-cost route. Subsequently, we generalize the algorithm to handle unknown target information, allowing the robot to scan and incrementally construct the target graph online while conducting robotic footprint coverage. The online polynomial-time SIFC planning algorithm minimizes the total robot traveling distance, guarantees complete sensing coverage of the entire workspace, and achieves near-optimal robotic footprint coverage, as demonstrated through empirical experiments. For the demonstrated application, we design coordinated nozzle motion control with the planned robot trajectory to efficiently fill all cracks within the robot’s footprint. Experimental results are presented to illustrate the algorithm’s design, performance, and comparisons. The SIFC algorithm offers a high-efficiency motion planning solution for various robotic applications requiring simultaneous sensing and actuation coverage.
@@ I Introduction
Surface cracks commonly exist in civil infrastructure, such as road and bridge deck surfaces, parking lots, airport runways, etc. To prevent crack growth and mitigate further deterioration, it is necessary to fill these cracks with appropriate materials in the early stages of crack appearance. It is labor-intensive, time-consuming, and expensive to repair the abundant existence of cracks in civil infrastructure by human workers. Robotics and automation technologies provide a promising tool to enable cost-effective, safe, and high-efficiency civil infrastructure maintenance. Mobile robot- or vehicle-based inspection systems were used for crack detection and maintenance on highways (e.g., [1, 2]). These systems were not fully automated, and human inspectors were still involved in operations. The developed autonomous robots in [3, 4, 5] mainly focused on nondestructive bridge deck inspection and crack mapping and repair were discussed. Crack mapping and filling can be viewed and generalized as a simultaneous robotic sensor-based inspection and footprint coverage (SIFC) problem. The onboard sensing system (e.g., camera) needs to detect all unknown targets (i.e., cracks) in the workspace, and at the same time, the robot must physically cover all the detected targets within its footprint to conduct the repair action. The complexity of the SIFC problem lies in simultaneously achieving the above-mentioned two complete coverage tasks of unknown targets using one robotic platform. The onboard target detection sensing range and the robot footprint have different sizes and geometric shapes. In addition, the robot footprint coverage task (e.g., filling cracks) might involve motion dynamics and control constraints, which is different from the coupled (passive) sensing coverage task. Although inspired by autonomous robotic crack mapping and filling, the SIFC is indeed a fundamental robot motion planning problem in other applications, such as sensing and cleaning dirty surfaces, finding and collecting mines, etc.
@@
The SIFC planning is related to the covering salesman problem, a traveling salesman problem variant in which an agent is required to travel the shortest distance to visit all specified neighborhoods in each city. However, without any knowledge of the locations of each city, the agent must inspect every point in the environment. Other variants include the art gallery problem and the watchman tour problem. However, these problems do not require the footprint coverage of the targets. All of these problems are NP-hard, and therefore, obtaining optimal solutions is only feasible for very limited problem domains [6]. To detect unknown targets, robotic exploration can be applied [7, 8]. In time-critical applications, most exploration paradigms are not efficient due to backtracking, where the robot might return to the same location more than once [9].
@@
Coverage path planning is widely used to explore the unknown environment by determining an optimal path that passes over all points of an area of interest while avoiding obstacles [10, 11]. Based on whether prior environmental information is known or not, coverage path planning methods can be classified into offline or online methods. In offline approaches, known information about the environment is used to produce the shortest or fastest path [12], while in online approaches, sensor information is used to plan coverage motion point-by-point. Many online coverage planning approaches use heuristics to navigate to the nearest cell [13], the cell with the highest potential [14, 15], or the cell with the lowest cost [16]. However, as they do not optimize the coverage path from a global perspective, these online planning approaches might fall into the local extremum. When the local extremum occurs, online approaches employ heuristic strategies such as backtracking procedures [17] or potential surfaces [14] to passively identify the next waypoint. These heuristic strategies prevent the robot from getting stuck in the local extremum; however, the robot visits the covered area repeatedly. Alternatively, the coverage path planning approaches in [18, 12] avoided the local extremum by using spanning trees, where the robot was guided to circumvent the virtual tree. These algorithms, however, did not account for partially occupied cells or special tree nodes and therefore resulted in incomplete coverage or repeated coverage [15].
@@
The Morse-based cellular decomposition (MCD) [19] and generalized Boustrophedon decomposition [20] guarantee the complete coverage of an unknown environment. The MCD methods ensure encountering all the critical points of the decomposition online [21, 22, 23]. Those algorithms perform an online decomposition such that the areas are covered completely by back-and-forth motions. However, no optimality was claimed for the planned paths. The work in [12] proposed an optimal coverage of a known environment. Grid-based approaches for planning a complete coverage path were also studied [24]. But those approaches restricted the space and robot motion to the grids and not to any arbitrary points in the workspace. Readers can further refer to [10, 25] for extensive surveys on coverage-planning strategies.
@@
Online re-planning of the robot’s path is needed when new information is obtained by the onboard sensors [26]. A graph-based Simplex method was presented to solve the re-planning problem [27]. The anytime dynamic A* algorithm is a generic graph-based re-planning scheme that generates bounded sub-optimal solutions when the map changes [28]. Navigation and coverage planning for autonomous underwater vehicles is a closely related problem (e.g., [29]). The existing work, however, focuses only on the unknown environment’s exploration task and does not consider additional simultaneously performing other tasks such as robot footprint coverage. Another piece of closely related work is the efficiently autonomous robotic cleaning or vacuuming [30, 31]. The learned dirt map was used for planning the robot to clean a set of cells. And this is different from the SIFC problem because the initial learning process is costly or infeasible for applications such as the online construction of the crack map for repair. The probabilistic planner in [32] used Monte Carlo localization to achieve complete coverage path planning, but it suffered from accumulating localization errors. Such probabilistic planners can perform online planning in dynamic environments, but they often have a higher computational complexity, rely on certain modeling assumptions (e.g., distributions of obstacles and uncertainty), and provide solutions that are only probabilistic optimal.
@@
In this paper, we present a set of motion planning algorithms for the SIFC with the application of robotic crack-filling in civil infrastructure. Instead of directly solving the SIFC problem, we first discuss the motion planning of a mobile robot to physically cover the target (i.e., cracks) for a given known target map; that is, no complete sensing coverage is considered. We decouple the two coverage planning tasks using classic cell decomposition methods to achieve complete sensing and robotic footprint coverage of the targets. A near-optimal complete footprint coverage plan is proposed to guide the robot, but with an offline-constructed target graph. The least-cost route is selected to traverse the constructed target graph. Finally, we propose a complete online motion planning solution for the SIFC problem, called the online sensor-based complete coverage (${\tt oSCC}$) algorithm. The ${\tt oSCC}$ detects unknown targets using onboard sensors when the robot traverses the targets to conduct the filling actuation at the shortest distance. In experiments, the crack-filling robot is equipped with four omni-directional wheels to perform arbitrary direction motion, and an $XY$-table mechanism is used to drive a fluid nozzle to conduct filling action. Motion control of omni-directional-wheel robots was reported in [33, 34]. We formulate motion coordination between the mobile robot and the nozzle movements into a nonlinear model predictive control (MPC). We conducted extensive experiments to validate and demonstrate the proposed planning and control algorithms.
@@
The main contribution of this work is the development of a new, complete, and empirically near-optimal motion planning and control approach for the SIFC problem in robotic crack mapping and filling applications. The proposed planning algorithms offer two key attractive features. Firstly, the novel ${\tt oSCC}$ algorithm guarantees complete sensing coverage of the free space in the entire workspace while simultaneously achieving complete robotic footprint coverage of the detected targets. This algorithm aims to minimize the total distance traveled by the robot and achieves a near-optimal path in polynomial time. Secondly, the motion control of the robotic filling mechanism is coordinated with the planned mobile robot trajectory, enabling efficient execution of the robot’s footprint task. The coupling between the sensor-based inspection and the onboard footprint coverage actuation is formulated and resolved by a new coordinated robot control design. Compared to the presented conference publications [35, 36], this paper introduces additional analyses for the motion planning and control design. It also presents extensive new experiments and detailed discussions using an upgraded crack-filling robot platform for improved motion planning and control. These developments provide valuable insights and a comprehensive understanding of the proposed approach in various scenarios, enhancing the applicability and reliability of the system.
@@
The rest of the paper is organized as follows: We discuss the problem statement and overview of the planning and control algorithms in Section II. Section III presents the basic footprint coverage planning with known target information. Section IV extends to sensor-based online coverage planning. Section V discusses the robotic control for crack-filling actuation. We present the experimental setup and results in Sections VI and VII, respectively. Finally, we summarize the concluding remarks in Section VIII.
@@ II-A Problem Statement
Fig. 1 shows an illustrative SIFC setup for the robotic crack-filling application. We consider the coverage planning and motion control for a robot, denoted as $\mathbf{R}$, which is equipped with a crack detection sensor and filling actuator. $\mathbf{R}$ is in a compact free workspace $\mathcal{W}\subset\mathbb{R}^{2}$. In general, the free space can be in any shape with a finite number of obstacles. For simplicity, we consider a known rectangular obstacle-free workspace with a size of $l\times w$, and the results are extendable to any other free spaces with different shapes. Readers can refer to [19, 21] for more details. Robot $\mathbf{R}$ needs to completely cover $\mathcal{W}$ to detect all unknown cracks (as targets) and simultaneously fill the detected cracks. $\mathbf{R}$ is assumed to freely move in any arbitrary direction in $\mathcal{W}$. The robot’s footprint, denoted as $\mathcal{F}$, is a circular area around its geometric center with a radius of $a$. Any cracks within $\mathcal{F}$ can be repaired by the filling actuator. The onboard target detection sensor (e.g., a panoramic camera) can identify any cracks within a circular area of a radius of $S$ around the robot’s geometric center. The target detection range is larger than the robot footprint size, that is, $S\geq a$.
@@
To focus on the coverage planning problem, robot $\mathbf{R}$ is assumed to know its location in $\mathcal{W}$. Crack widths are assumed constant and $\mathbf{R}$ only needs to consider the crack length as footprint coverage. We also assume that at any time the nozzle-filling motion and the robot motion are coordinated so that any targets under the robot’s footprint can be reached by the nozzle. Therefore, we only consider the minimization of the robot’s total traveling distance as the goal of the planner.
@@ II-B Algorithm Design Overview
To solve the above SIFC problem, we present a set of algorithmic developments. Fig. 2 illustrates an overview of the planning algorithms and their relationships. First, we describe a robotic coverage planning algorithm, called graph-based coverage (${\tt GCC}$), to drive robot $\mathbf{R}$ to cover all targets by its footprint under a given target map. Note that ${\tt GCC}$ does not cover the entire workspace $\mathcal{W}$ by using onboard detection sensors, but it serves as a basic planning module to cover the given target map by robotic footprint $\mathcal{F}$.
@@
We then present a sensor-based complete coverage of workspace $\mathcal{W}$ assuming known target maps, which is denoted as the $\tt SCC$ algorithm. The $\tt SCC$ algorithm initially constructs a graph map from the provided target maps and subsequently utilizes the $\tt GCC$ to explore the entire workspace $\mathcal{W}$ and guide robot $\bf{R}$ to achieve footprint coverage of all targets while minimizing travel distance. Serving as the foundation for addressing the SIFC problem, the $\tt SCC$ algorithm plays a crucial role. By leveraging the insights gained from the edge connections observed in the $\tt SCC$ algorithm and its properties of completeness and near-optimality, we extend this approach to handle scenarios with unknown target information, such as unknown cracks. This extension leads to the development of an online $\tt SCC$ algorithm, denoted as $\tt oSCC$, which is derived from $\tt SCC$ by relaxing the assumption of given target maps to achieve near-optimal sensor-based target detection and real-time robot footprint coverage to completely cover $\mathcal{W}$. Finally, crack filling control is designed to drive the onboard actuation mechanism to fill the detected targets with coordinated robot motion given by $\tt oSCC$. Both $\tt SCC$ and $\tt oSCC$ use ${\tt GCC}$ as part of the algorithmic module. The crack-filling control has to follow the dynamic constraints of the actuator and mechanical systems with the coupled robot motion. The sensor-based detection coverage by the $\tt oSCC$ algorithm and robotic footprint coverage design together provide the SIFC solution. In the next two sections, we present the $\tt GCC$, $\tt SCC$, and $\tt oSCC$ algorithms in detail.

### v2

@@ Abstract
We present a simultaneous sensor-based inspection and footprint coverage (SIFC) planning and control design with applications to autonomous robotic crack mapping and filling. The main challenge of the SIFC problem lies in the coupling of complete sensing (for mapping) and robotic footprint (for filling) coverage tasks. Initially, we assume known target information (e.g., cracks) and employ classic cell decomposition methods to achieve complete sensing coverage of the workspace and complete robotic footprint coverage using the least-cost route. Subsequently, we generalize the algorithm to handle unknown target information, allowing the robot to scan and incrementally construct the target map online while conducting robotic footprint coverage. The online polynomial-time SIFC planning algorithm minimizes the total robot traveling distance, guarantees complete sensing coverage of the entire workspace, and achieves near-optimal robotic footprint coverage, as demonstrated through experiments. For the demonstrated application, we design coordinated nozzle motion control with the planned robot trajectory to efficiently fill all cracks within the robot’s footprint. Experimental results illustrate the algorithm’s design, performance, and comparisons. The SIFC algorithm offers a high-efficiency motion planning solution for various robotic applications requiring simultaneous sensing and actuation coverage.
@@
©2024 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission. See https://www.ieee.org/publications/rights/index.html for more information. This article has been accepted for publication in IEEE Transactions on Robotics. Citation information: DOI 10.1109/TRO.2024.3392077.
@@ I Introduction
Surface cracks commonly exist in civil infrastructure, such as road and bridge deck surfaces, parking lots, airport runways, etc. To prevent crack growth and mitigate further deterioration, it is necessary to fill these cracks with appropriate materials in the early stages of crack appearance. Repairing the abundant existence of cracks in civil infrastructure by human workers is labor-intensive, time-consuming, and expensive. Robotics and automation technologies provide a promising tool to enable cost-effective, safe, and high-efficiency civil infrastructure maintenance. Mobile robot- or vehicle-based inspection systems were used for crack detection and maintenance on highways (e.g., [1, 2]). However, these systems were not fully automated, and human inspectors were still involved in operations. The autonomous robots developed in [3, 4, 5] mainly focused on nondestructive bridge deck inspection, and although crack mapping and repair were discussed, they were not the primary focus. Crack mapping and filling can be viewed and generalized as a simultaneous robotic sensor-based inspection and footprint coverage (SIFC) problem. The onboard sensing system (e.g., camera) needs to detect all unknown targets (i.e., cracks) in the workspace, and at the same time, the robot must physically cover all the detected targets within its footprint to conduct the repair action. The complexity of the SIFC problem lies in simultaneously achieving the above-mentioned two complete coverage tasks of unknown targets using one robotic platform. The onboard target detection sensing range and the robot footprint have different sizes and geometric shapes. Additionally, the robot footprint coverage task (e.g., filling cracks) might involve motion dynamics and control constraints, which are different from the coupled (passive) sensing coverage task. Although inspired by autonomous robotic crack mapping and filling, the SIFC is indeed a fundamental robot motion planning problem in other applications, such as sensing and cleaning dirty surfaces, finding and collecting mines, etc.
@@
The SIFC planning is related to the covering salesman problem, a variant of the traveling salesman problem where an agent must travel the shortest distance to visit all specified neighborhoods in each city. However, unlike traditional scenarios where the agent has prior knowledge of each city’s location, in the SIFC problem, the agent must inspect every point in the environment to detect unknown targets. Other related problems include the art gallery problem and the watchman tour problem, but these do not involve the footprint coverage of targets. It’s worth noting that all these problems are NP-hard, making obtaining optimal solutions feasible only for very limited problem domains [6]. Robotic exploration techniques can be used to detect unknown targets [7, 8]. However, in time-critical applications, many exploration paradigms suffer from inefficiencies due to backtracking, where the robot may revisit the same location more than once [9].
@@
Coverage path planning explores environments by determining an optimal path that covers all points of interest while avoiding obstacles [10, 11]. These methods are classified into offline or online methods based on whether prior environmental information is known. In offline approaches, known information about the environment is used to produce the shortest or fastest path [12], while in online approaches, sensor information is used to plan coverage motion point-by-point. Many online strategies used heuristics to navigate to the nearest cell [13], the cell with the highest potential [14, 15], or the cell with the lowest cost [16]. However, as they did not optimize the coverage path from a global perspective, these online planning approaches might fall into local extrema. To prevent the robot from getting stuck in local extrema, online approaches employed heuristic strategies such as backtracking procedures [17] or potential surfaces [14] to passively identify the next waypoint. These heuristic strategies prevent the robot from getting stuck in local extrema; however, they may cause the robot to repeatedly visit the covered area. Alternatively, coverage path planning approaches in [18, 12] avoided local extrema by using spanning trees, guiding the robot to circumvent the virtual tree. However, these algorithms did not account for partially occupied cells or special tree nodes, resulting in incomplete or repeated coverage [15].
@@
Morse-based cellular decomposition (MCD) [19] and generalized Boustrophedon decomposition [20] guarantee complete coverage of an unknown environment. MCD methods ensure encountering all critical points of the decomposition online [21, 22, 23]. These algorithms perform online decomposition such that the areas are covered completely by back-and-forth motions. However, no optimality is claimed for the planned paths. The work in [12] proposed an optimal coverage of a known environment. Grid-based approaches for planning a complete coverage path were also studied [24]. However, these approaches restrict space and robot motion to grids rather than arbitrary points in the workspace. Readers can refer to [10, 25] for extensive surveys on coverage planning strategies.
@@
Online re-planning of the robot’s path is necessary when new information is obtained by onboard sensors [26]. A graph-based Simplex method was presented to solve the re-planning problem [27]. The anytime dynamic A* algorithm is a generic graph-based re-planning scheme that generates bounded sub-optimal solutions when the map changes [28]. Navigation and coverage planning for autonomous underwater vehicles is a closely related problem (e.g., [29]). However, existing work focuses solely on the exploration task of unknown environments and does not consider simultaneously performing other tasks such as robot footprint coverage. Another closely related piece of work is efficient autonomous robotic cleaning or vacuuming [30, 31], where a learned dirt map is used for planning the robot to clean a set of cells. This differs from the SIFC problem because the initial learning process may be costly or infeasible for applications such as the online construction of a crack map for repair. The probabilistic planner in [32] utilized Monte Carlo localization for complete coverage path planning but suffered from accumulating localization errors. While such probabilistic planners can perform online planning in dynamic environments, they often entail higher computational complexity, rely on specific modeling assumptions (e.g., distributions of obstacles and uncertainty), and provide solutions that are only probabilistically optimal.
@@
In this paper, we present a set of motion planning algorithms for the SIFC problem with the application of robotic crack-filling in civil infrastructure. Instead of directly solving the SIFC problem, we first discuss the motion planning of a mobile robot to physically cover the targets (i.e., cracks) for a given known target map; that is, no complete sensing coverage is considered. We decouple the two coverage planning tasks using classic cell decomposition methods to achieve complete sensing and robotic footprint coverage of the targets. A near-optimal complete footprint coverage plan is proposed to guide the robot, but with an offline-constructed target graph. The least-cost route is selected to traverse the constructed target graph. Finally, we propose a complete online motion planning solution for the SIFC problem, called the online sensor-based complete coverage (${\tt oSCC}$) algorithm. The ${\tt oSCC}$ detects unknown targets using onboard sensors when the robot traverses the targets to conduct the filling actuation at the shortest distance. In experiments, the crack-filling robot is equipped with four omni-directional wheels to perform arbitrary direction motion, and an $XY$-table mechanism is used to drive a fluid nozzle for the filling action. Motion control of omni-directional-wheel robots was reported in [33, 34]. We formulate motion coordination between the mobile robot and the nozzle movements into a nonlinear model predictive control (MPC). Extensive experiments validate and demonstrate the proposed planning and control algorithms.
@@
The main contribution of this work is the development of a new, complete, and empirically near-optimal motion planning and control approach for the SIFC problem in robotic crack mapping and filling applications. The proposed planning algorithms offer two key attractive features. First, the novel ${\tt oSCC}$ algorithm guarantees complete sensing coverage of the free space in the entire workspace while simultaneously achieving complete robotic footprint coverage of the detected targets. This algorithm aims to minimize the total distance travelled by the robot and achieves a near-optimal path in polynomial time. Second, the motion control of the robotic filling mechanism is coordinated with the planned mobile robot trajectory, enabling efficient execution of the robot’s footprint task. The coupling between the sensor-based inspection and the onboard footprint coverage actuation is formulated and resolved by a new coordinated robot control design. Compared to the presented conference publications [35, 36], this paper introduces additional analyses for the motion planning and control design. It also presents extensive new experiments and detailed discussions using an upgraded crack-filling robot platform for improved motion planning and control. These developments provide valuable insights and a comprehensive understanding of the proposed approach in various scenarios, enhancing the applicability and reliability of the system.
@@
The rest of the paper is organized as follows. In Section II, we outline the problem statement and provide an overview of the planning and control algorithms. Section III discusses basic footprint coverage planning with known target information, while Section IV extends to sensor-based online coverage planning. Robotic control for crack-filling actuation is covered in Section V. Experimental setup and results are presented in Sections VI and VII, respectively. Finally, concluding remarks are summarized in Section VIII.
@@ II-A Problem Statement
Fig. 1 shows an illustrative SIFC setup for the robotic crack-filling application. We consider the coverage planning and motion control for a robot, denoted as $\mathbf{R}$, which is equipped with a crack detection sensor and filling actuator. $\mathbf{R}$ is in a compact free workspace $\mathcal{W}\subset\mathbb{R}^{2}$. In general, the free space can be in any shape with a finite number of obstacles. For simplicity, we consider a known rectangular obstacle-free workspace with a size of $l\times w$, and the results are extendable to any other free space with different shape (readers can refer to [19, 21] for more details). Robot $\mathbf{R}$ needs to completely cover $\mathcal{W}$ to detect all unknown cracks (as targets) and simultaneously fill the detected cracks. $\mathbf{R}$ is assumed to freely move in any arbitrary direction in $\mathcal{W}$. The robot’s footprint, denoted as $\mathcal{F}$, is a circular area around its geometric center with a radius of $a$. Any cracks within $\mathcal{F}$ can be repaired by the filling actuator. The onboard target detection sensor (e.g., a panoramic camera) can identify any cracks within a circular area of a radius of $S$ around the robot’s geometric center. The target detection range is larger than the robot footprint size, that is, $S\geq a$.
To focus on the coverage planning problem, robot $\mathbf{R}$ is assumed to know its location in $\mathcal{W}$. Crack widths are assumed to be constant, and $\mathbf{R}$ only needs to consider the crack length for footprint coverage. Additionally, we assume that the nozzle-filling motion and the robot’s motion are coordinated at all times, ensuring that any targets under the robot’s footprint can be reached by the nozzle. Therefore, our primary goal for the planner is to minimize the robot’s total traveling distance.
@@ II-B Algorithm Design Overview
To solve the above SIFC problem, we present a set of algorithmic developments. Fig. 2 illustrates an overview of the planning algorithms and their relationships. First, we describe a robotic coverage planning algorithm, called graph-based coverage (${\tt GCC}$), to drive robot $\mathbf{R}$ to cover all targets by its footprint under a given target map. Note that ${\tt GCC}$ does not cover the entire workspace $\mathcal{W}$ using onboard detection sensors, but it serves as a basic planning module to cover the given target map by robotic footprint $\mathcal{F}$.
@@
We then present a sensor-based complete coverage of workspace $\mathcal{W}$ assuming known target maps, which is denoted as the $\tt SCC$ algorithm. The $\tt SCC$ algorithm initially constructs a graph map from the provided target maps and subsequently utilizes the $\tt GCC$ to explore the entire workspace $\mathcal{W}$ and guide robot $\bf{R}$ to achieve footprint coverage of all targets while minimizing travel distance. Serving as the foundation for addressing the SIFC problem, the $\tt SCC$ algorithm plays a crucial role. By leveraging the insights gained from the edge connections observed in the $\tt SCC$ algorithm and its properties of completeness and near-optimality, we extend this approach to handle scenarios with unknown target information. This extension leads to the development of an online $\tt SCC$ algorithm, denoted as $\tt oSCC$, which is derived from $\tt SCC$ by relaxing the assumption of given target maps to achieve near-optimal sensor-based target detection and real-time robot footprint coverage to completely cover $\mathcal{W}$. Finally, crack filling control is designed to drive the onboard actuation mechanism to fill the detected targets with coordinated robot motion given by $\tt oSCC$. Both $\tt SCC$ and $\tt oSCC$ use ${\tt GCC}$ as part of the algorithmic module. The crack-filling control has to follow the dynamic constraints of the actuator and mechanical systems, considering the coupled robot motion. The sensor-based detection coverage offered by the $\tt oSCC$ algorithm and the robotic footprint coverage design collectively provide the SIFC solution. In the next two sections, we present the $\tt GCC$, $\tt SCC$, and $\tt oSCC$ algorithms in detail.